# biblio.bib

### v2

@@ 4.1 Technical claims and proofs
Observe that the improvement of the entire sequence of $\beta$-improving sets $((A_{i},\ell_{i}))_{1\leq i\leq h}$ is at least $\beta w(V_{H})$, due to the definition of sequence of $\beta$-improving sets Definition 3.3. Now since the maximum improvement in the potential of any labeling in $G$ is at most $\Lambda\cdot n$, the overall improvement of the sequence itself cannot be larger than $\Lambda\cdot n$. This implies that $\beta w(V_{H})\leq\Lambda\cdot n$, which, in turn, implies $w(V_{H})\in O(\Lambda n/\beta)$.
Our ball-growing argument will start with a node $v=A_{j}\in V_{H}$. We then grow the ball by one additional hop at every step: our “candidate” sets are $C_{0}=\{A_{j}\},C_{1}=\mathcal{N}_{1}[\{A_{j}\}],\ldots,C_{i}=\mathcal{N}_{i}[\{A_{j}\}]$, where these sets of vertices are subsets of $V_{H}$. We stop the process when $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1}\setminus C_{i})=w(C_{i+1})-w(C_{i})$, for $K=(\beta-\varepsilon)+\varepsilon\Delta^{r}\Lambda$. Notice that there must be an index $i=O(K\operatorname{log}(\Lambda n/\beta)/\varepsilon)$ for which the statement holds. By contradiction, suppose that the statement does not hold for all $i\leq\left\lceil KM\operatorname{log}(\Lambda n/\beta)\right\rceil$, where $M$ is a large enough constant. Then, we have that $w(C_{i+1})-w(C_{i})\geq\frac{\varepsilon}{K}w(C_{i})$ for all $i<KM\operatorname{log}(\Lambda n/\beta)/\varepsilon$. Let $j=\left\lceil KM\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Then,
$w(C_{j})\geq(1+\varepsilon/K)w(C_{j-1})\geq(1+\varepsilon/K)^{KM\operatorname{log}(\Lambda n/\beta)/\varepsilon}.$
Again, by exploiting that $(1+x/y)^{y}\geq\exp[xy/(x+y)]$ for all positive $x,y$, we get that
$w(C_{j})\geq\exp\left[\frac{KM\operatorname{log}(\Lambda n/\beta)}{\varepsilon+K}\right]\geq(\Lambda n/\beta)^{M/2}.$
If $M\geq 3$, then we get a contradiction since $w(C_{j})\leq w(V_{H})\leq\Lambda n/\beta$.
Let $i^{\star}=O(K\operatorname{log}(\Lambda n/\beta)/\varepsilon)$ be the first index that realizes $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1})-w(C_{i})$. We now argue that, in $G$, the pair $(A,\ell^{\star}_{A})$, with
$A=\bigcup_{A_{j}\in C_{i^{\star}+1}}A_{j},$
is an improving set with $\operatorname{IR}(A,\ell,\ell^{\star}_{A})\geq\beta-\varepsilon$. Let
$A^{\prime}=\bigcup_{A_{j}\in C_{i^{\star}}}A_{j},$
@@
and let $s$ be the maximum index such that $A_{s}$ is in $C_{i^{\star}+1}$. Denote by $\ell_{A}$ the labeling obtained by setting $\ell_{A}=\ell_{s}$ for all nodes in $A$ and their incident half-edges, and $\ell_{A}=\ell$ for all other nodes and half-edges.
We argue that for all of the nodes close to $A^{\prime}$ the labeling $\ell_{A}$ looks the same as applying the entire improving sequence up until labeling $\ell_{s}$. The only nodes that are able to see a difference are the ones close to $A\setminus A^{\prime}$.
To this end notice that we can obtain $\ell_{s}$ by essentially just applying all of the changes of $\ell_{1},\ell_{2},\ldots,\ell_{s}$ sequentially one after the other. Here, each $\ell_{i}$ acts only on $A_{i}$ by the definition of improving sets $(A_{i},\ell_{i})$. So if we define $\ell_{0}=\ell$ and $A_{S}=\bigcup_{1\leq i\leq s}A_{i}$, we get
$\operatorname{Imp}(A_{s},\ell,\ell_{s})=\sum_{1\leq i\leq s}\operatorname{Imp}(A_{i},\ell_{i-1},\ell_{i})$
We want to achieve something similar, but for $\ell_{A}$, so we need to only apply the changes for the $A_{i}\subseteq A$. Note that the $A_{i}\subseteq A$ are essentially an arbitrary subsequence of our original sequence. So it is a priori unclear how the result of applying these changes in sequence will look like (while potentially skipping crucial steps).
@@
Furthermore, let $A_{i_{1}},\ldots,A_{i_{t}}$ be the elements of $C_{i^{\star}+1}$, with $i_{1}<\ldots<i_{t}=s$. Let $\ell_{i_{1}},\ell_{j_{2}},\ldots,\allowbreak\ell_{i_{t}}=\ell_{s}$, be the corresponding labelings. We further define $f_{i_{0}}=\ell$ and then our restricted labelings $f_{i_{j}}$ as the labeling that outputs $\ell_{{}_{j}}$ on all of $A_{i_{j}}$, but outputs $f_{i_{j-1}}$ everywhere else. Now we can obtain $\ell_{A}$, by sequentially applying all of the changes $f_{i_{1}},f_{i_{2}},\ldots,f_{i_{t}}=\ell_{A}$.
Importantly, for any node in $A^{\prime}$, the sequence of labelings $f_{i_{1}},f_{i_{2}},\ldots,f_{i_{t}}=\ell_{A}$ is indistinguishable from the full sequence $\ell_{1},\ell_{2},\ldots,\ell_{s}$. This is true, because the sequence $f_{i_{1}},f_{i_{2}},\ldots,f_{i_{t}}$ contains all of the changes to nodes at distance at most $r$ from any node in $A^{\prime}$. (Notice that if a node $v$ in the $r$ hop neighborhood of $A^{\prime}$ can witness any change of a labeling $\ell_{x}$, then $v$ must also be at distance at most $r$ from $A_{x}$, so $\operatorname{dist}(A^{\prime},A_{x})\leq 2r$ and so $A_{x}$ must be adjacent to $A^{\prime}$ and so $A_{x}\in A$)
As a result
$\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})=\operatorname{Imp}(A_{i_{j}},\ell_{i_{j}-1},\ell_{i_{j}})\geq\beta\cdot w(A_{i_{j}})$
if $A_{i_{j}}\in A^{\prime}$. On the other hand, if $A_{i_{j}}\in A\setminus A^{\prime}$ things could be bad, but the increase in potential is at most $\Delta^{r}\Lambda\cdot w(A_{i_{j}})$, since the changes in $A_{j_{i}}$ can influence at most the $r$ hop neighborhood and each node can decrease by at most $\Lambda$. We obtain the following
$\operatorname{Imp}(A,\ell,\ell_{A})=$ $\sum_{j=1}^{s}\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})$
=\ = $\sum_{A_{i_{j}}\in A^{\prime}}\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})+\sum_{A_{i_{j}}\in A\setminus A^{\prime}}\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})$
=\ = $\sum_{A_{i_{j}}\in A^{\prime}}\operatorname{Imp}(A_{i_{j}},\ell_{i_{j}-1},\ell_{i_{j}})-\sum_{A_{i_{j}}\in A\setminus A^{\prime}}\Delta^{r}\cdot\Lambda\cdot w(A_{i_{j}})$
=\ = $\beta w(C_{i^{\star}})-\Delta^{r}\cdot\Lambda\cdot(w(C_{i^{\star}}+1)-w(C_{i^{\star}}))$
@@
=\ = $w(C_{i^{\star}})\left(\beta-\frac{\varepsilon\Delta^{r}\Lambda}{K}\right).$
It is then sufficient to set $K$ large enough so that
$w(C_{i^{\star}})\left(\beta-\frac{\varepsilon\Delta^{r}\Lambda}{K}\right)\geq w(C_{i^{\star}+1})\left(\beta-\varepsilon\right).$
Note that $w(C_{i^{\star}+1})\leq w(C_{i^{\star}})(1+\varepsilon/K)$. Hence, we choose $K$ such that
$w(C_{i^{\star}})\left(\beta-\frac{\varepsilon\Delta^{r}\Lambda}{K}\right)\geq w(C_{i^{\star}})\left(1+\frac{\varepsilon}{K}\right)\left(\beta-\varepsilon\right).$
Choosing any $K\geq(\beta-\varepsilon)+\varepsilon\Delta^{r}\Lambda$ does the job (and hence $K=(\beta-\varepsilon)+\varepsilon\Delta^{r}\Lambda$) is sufficient.
We finally argue that $A$ is fully contained in the radius-$O(K\operatorname{log}(\Lambda n/\beta)\operatorname{log}n/\varepsilon^{2})$ neighborhood of $A_{j}$ (in $G$) and, in particular, since $A_{j}$ has strong diameter at most $O(\operatorname{log}n/\varepsilon)$, $A$ lies in the radius-$O(K\operatorname{log}(\Lambda n/\beta)\operatorname{log}n/\varepsilon^{2})$ neighborhood of any node of $A_{j}$. Since all $A_{i}$s have strong diameter at most $O(\operatorname{log}n)$, and $C_{i^{\star}+1}$ has diameter $i^{\star}+1=O(K\operatorname{log}(\Lambda n/\beta)/\varepsilon)$ in $H$, then nodes of $G$ can “simulate” the supergraph $H$ and “find” $C_{i^{\star}+1}$ with locality $O(K\operatorname{log}(\Lambda n/\beta)\operatorname{log}n/\varepsilon^{2})$.
By noticing that $\ell^{\star}_{A}$ can only increase the improving ratio w.r.t. $\ell_{A}$, we conclude the proof of Lemma 4.8. ∎
@@
We first apply Lemma 4.8, and we obtain an improving set $(A,\ell_{A})$ of diameter $c_{3}\operatorname{log}^{2}/\varepsilon^{2}$ with the desired properties. Then, we take the minimal version of it, that is, $(A^{\prime},\ell^{\star}_{A^{\prime}})$, with $A^{\prime}\subseteq A$, and we apply Lemma 4.5, which yields the thesis. ∎

### v3

@@ 4.1 Technical claims and proofs
Observe that the improvement of the entire sequence of $\beta$-improving sets $((A_{i},\ell_{i}))_{1\leq i\leq h}$ is at least $\beta\cdot w(V_{H})$, due to the definition of sequence of $\beta$-improving sets. Now since the maximum improvement in the potential of any labeling in $G$ is at most $\Lambda\cdot n$, the overall improvement of the sequence itself cannot be larger than $\Lambda\cdot n$. This implies that $w(V_{H})\leq\Lambda\cdot n/\beta$.
Fix any $j\in[h]$, and consider an arbitrary $v\in A_{j}$. Let $x=\sigma(A_{j})\in V_{H}$. We define the “candidate sets” with regard to $z$ as $C_{i}=\mathcal{N}_{i}[z]$ for $i\geq 0$. Let $k$ be the minimum integer $i$ such that $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1}\setminus C_{i})=w(C_{i+1})-w(C_{i})$, for $K=(\beta-\varepsilon)+(\Delta^{r}+1)\Lambda$. We claim that $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. By contradiction, suppose that $k>\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Let $j=\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Then, we have that $w(C_{i+1})-w(C_{i})\geq\frac{\varepsilon}{K}w(C_{i})$ for all $i\leq j$. Then,
$w(C_{j})\geq(1+\varepsilon/K)w(C_{j-1})\geq(1+\varepsilon/K)^{3K\operatorname{log}(\Lambda n/\beta)/\varepsilon}.$
Again, by exploiting that $(1+x/y)^{y}\geq\exp[xy/(x+y)]$ for all positive $x,y$, and that $\varepsilon<\lambda\leq\Lambda$, we get that
$w(C_{j})\geq\exp\left[\frac{KM\operatorname{log}(\Lambda n/\beta)}{\varepsilon+K}\right]\geq(\Lambda n/\beta)^{3/2}.$
This is a contradiction since $w(C_{j})\leq w(V_{H})\leq\Lambda n/\beta$.
Let
$A$ $=\{i\in[h]\mid A_{i}=\sigma^{-1}(x)\text{ for some }x\in C_{k+1}\},\text{ and}$
$A^{\prime}$ $=\{i\in[h]\mid A_{i}=\sigma^{-1}(x)\text{ for some }x\in C_{k}\}.$
@@
Furthermore, let $s=\max(A)$. We also define $\hat{A}=\cup_{i\in A}A_{i}$, and $\hat{A}^{\prime}=\cup_{i\in A^{\prime}}A_{i}$. Denote by $\ell_{A}$ the labeling obtained by setting $\ell_{A}=\ell_{s}$ for all nodes in $\hat{A}$ and their incident half-edges, and $\ell_{A}=\ell$ for all other nodes and half-edges.
@@
Let $i_{1},\ldots,i_{t}$ denote the elements of $A$, where $i_{1}<\ldots<i_{t}=s$. We define $f_{i_{0}}=\ell$ and then a new relabeling $f_{i_{j}}$ as the labeling that outputs $\ell_{i_{j}}$ on all the nodes in $A_{i_{j}}$ and their incident half-edges, but outputs $f_{i_{j-1}}$ everywhere else. Now we can obtain $\ell_{A}$ by sequentially applying the relabelings $f_{i_{1}},f_{i_{2}},\ldots,f_{i_{t}}=\ell_{A}$.
Importantly, for each $j\in[t]$, by the definition of $f_{i_{j}}$ and $E_{H}$, the labels of all nodes in $\mathcal{N}_{2r}[\hat{A}^{\prime}]$ and their incident half-edges given by $f_{i_{j}}$ are exactly those given by $\ell_{i_{j}}$: in fact, if the labeling of a node $u\in\mathcal{N}_{2r}[\hat{A}^{\prime}]$ (or its incident half-edges) is modified by $\ell_{i_{j}}$, then $u\in A_{i_{j}}$ and $\operatorname{dist}_{G}(A_{i_{j}},\hat{A}^{\prime})\leq 2r$, implying that $i_{j}\in A$.
As a result, if $i_{j}\in A^{\prime}$, then
$|.\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})=\operatorname{Imp}(A_{i_{j}},\ell_{i_{j-1}},\ell_{i_{j}})\geq\operatorname{Imp}(A_{i_{j}},\ell_{i_{j}-1},\ell_{i_{j}})\geq\beta\cdot\absolutevalue{A_{i_{j}}}$.
@@
$\operatorname{Imp}(\hat{A},\ell,\ell_{A})=\$$\sum_{j=1}^{t}\operatorname{Imp}(A_{i_{j}},f_{i_{j-1}},f_{i_{j}})$
[…]
We now show that, by the definition of $K$, it holds that
$w(C_{k})\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)\geq w(C_{k+1})\left(\beta-\varepsilon\right).$
Recall that $w(C_{k+1})\leq w(C_{k})(1+\varepsilon/K)$. Hence,
$w(C_{k})\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)$ $\geq w(C_{k+1})\cdot\frac{K}{K+\varepsilon}\cdot\left(\beta-\frac{\varepsilon(\Delta^{r}+1)\Lambda}{K}\right)$
$=w(C_{k+1})\cdot\frac{\beta K-\varepsilon(K-(\beta-\varepsilon))}{K+\varepsilon}$
$=w(C_{k+1})\cdot\frac{K(\beta-\varepsilon)+\varepsilon(\beta-\varepsilon)}{K+\varepsilon}$
$=w(C_{k+1})\cdot\left(\beta-\varepsilon\right)\cdot\frac{K+\varepsilon}{K+\varepsilon}$
$=w(C_{k+1})\cdot\left(\beta-\varepsilon\right).$
Note that $C_{k+1}=\mathcal{N}_{k+1}[z]$. Thus, the weak diameter of $\hat{A}$ in $G$ is at most $(2k+1)\cdot(2r+c_{3}\operatorname{log}n/\varepsilon)$. In particular, since $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$ and $K\leq\beta+(\Delta^{r}+1)\Lambda$, we have that the weak diameter of $\hat{A}$ in $G$ is at most $100((\beta+\Delta^{r}\Lambda)\cdot\operatorname{log}(\Lambda n/\beta)/\varepsilon+1)(r+c_{3}\operatorname{log}n/\varepsilon)$. Since by assumption $\Delta$ and $r$ are constants, we have that $\hat{A}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$ for some large enough constant $c_{2}>0$. By noticing that $\ell^{\star}_{\hat{A}}$ can only increase the improving ratio w.r.t. $\ell_{{A}}$, we conclude the proof of Lemma 4.8. ∎
@@
We first apply Lemma 4.8, and we obtain an improving set $(A,\ell_{A})$ with improving ratio at least $\beta-\varepsilon$ that is fully contained in $\mathcal{N}_{c_{2}\operatorname{log}^{2}/\varepsilon^{2}}[v]\cap(\cup_{i\in[h]}A_{i})$. Then, we take the minimal version of it, that is, $(A^{\prime},\ell^{\star}_{A^{\prime}})$, with $A^{\prime}\subseteq A$, and we apply Lemma 4.5, which yields the thesis. ∎